# 2]organization=European Synchrotron Radiation Facility, addressline=71 Avenue des Martyrs, city=Grenoble, postcode=F-38000, state=France, country=

### v1

@@ Abstract
keywords:
@@
Introduction
Intense narrow-band multi-MeV photon beams have a large potential in fields of basic and nuclear research, medicine and technical applications [1, 2, 3, 4]. Having in mind in particular photonuclear reactions for application purposes, the production mechanisms of such beams should be targeted, from an economical point of view, on electron accelerators with moderate energies in the GeV range or below. Bremsstrahlung radiation produced by an electron beam incident on a high-Z target has been widely used for many decades. Alternatively, also channeling radiation has been considered as an intense radiation source as well [5]. However, both sources feature broad band radiation characteristics. Facilities based on laser Compton scattering are already operational or currently under construction [3, Table 10] by which nearly mono-energetic $\gamma$-ray beams of high intensity can be produced.
Figure 1: Effect of a periodic variation of the lattice constant. A superlattice (undulator) crystal originates as shown by the red colored full line for one period. The effect of the tetragonal distortion of the equilibrium cubic unit cell has been described by M.B.H. Breese [6]. The electron beam enters in $z$ direction, called forward direction, which coincides with the [110] direction of the plane diamond backing (host) crystal. The $z^{\prime}$ direction is called undulator direction. The two coordinate systems $(x,z)$ and the $(x^{\prime},z^{\prime})$ are interconnected by a rotation transform with an angle $\delta$. Planes are formed perpendicular to the plane of drawing. Figure taken from [7].
Figure 2: Transverse potential landscape of a boron doped (110) diamond superlattice. Shown is one period with length of 5.0 $\mu$m, centrifugal potential omitted. A full sinusoidal period starts at $x=z$ = 0. Notice, the longitudinal $z$ coordinate is squeezed by four orders of magnitudes compared with the transverse $x$ one. It is shown in Methods Monte-Carlo simulation calculations that the barrier heights are modified as functions of $z$ by centrifugal forces. The quantity $\overline{x}$ is defined by Eq. (4).
In this work an alternative route towards narrow band photon sources was elaborated which is based on the channeling process of ultra-relativistic electrons in periodical bent crystals. The scheme resembles magnetic undulators, however, with period lengths in the multi-$\mu$m range rather than on a centimeter scale. Such undulator-like radiation sources in the MeV range and beyond have been comprehensively discussed by Korol and Solov’yov, see e.g. [8] and references cited therein.
The experimental challenge is to fabricate crystalline undulators with such small period lengths. First attempts were undertaken with strained layer Si1-xGex undulator crystals grown at the Aarhus University [9]. Of greater interest are undulators based on low $Z$ = 6 diamond, because of its radiation hardness, which can be produced by Chemical Vapour Deposition (CVD) with periodically varying concentration of boron. It is based on the functional dependence of the lattice constant on the boron density, e.g. [10, Fig. 1]. As schematically shown by Fig. 1, oscillating (110) planes result with a potential landscape depicted in Fig. 2.
An electron entering the crystal at the coordinates $x=z$ = 0 nearly parallel to the $z$ axis may be captured in the potential pocket with a barrier hight of 24 eV, and performs oscillations, thereby following the bent sinusoidal shape of the (110) plane. By means of the accompanied acceleration channeling radiation and undulator-like radiation will be emitted, respectively. The radiation is strongly forward directed, for undulator-like radiation typically in the order of 0.5 mrad for 1 GeV electrons. The emitted photon energy is in the $\gamma$-ray domain, for a period length of a few $\mu$m typically in the MeV range. It scales with the electron beam energy squared and inversely with the undulator period length. The capture probability and in turn the intensity of the $\gamma$ radiation requires high quality electron beams with a low angular spread. The reason is that the transverse energy $pv/2\cdot\vartheta_{x}^{2}$ is the relevant quantity for the dynamics of the electron in a potential landscape as shown in Fig. 2. The quantity $pv$ is in essence the electron beam energy and $\vartheta_{x}$ the angle of the electron with respect to the $z$ axis. Collisions of the beam electron with atomic nuclei and electrons comprising the crystal, potentially result in an increase of the transverse energy and de-channeling processes. That is another reason why the low $Z$ = 6 material of diamond is a good choice.
An experiment was performed by us at planar (110) channeling of electrons in a four-period diamond undulator with a period length $\lambda$ = 5.0 $\mu$m. Of paramount importance was the successful fabrication by the CVD process and diagnostics of such a boron doped superlattice. The experiment has been conducted with the high quality 855 MeV electron beam of the Mainz Microtron MAMI accelerator facility. Of crucial importance for the successful peak search at the predicted energy of 1.33 MeV was a strategy developed prior to the experiment. To this end Monte-Carlo simulation calculations were performed with a computer code developed by us, which facilitated the experimental search with optimized parameter predictions.
Our computer code fairly well reproduces the basic measured characteristics of the quasi-monochromatic $\gamma$ radiation as the peak energy, its width and the intensity. This fact encouraged us to predict, as an example, features of intense quasi-monochromatic 14.3 MeV radiation from an eight-period diamond undulator with $\lambda$ = 5 $\mu$m and a 3 GeV electron beam. Such a $\gamma$-ray source may be of potential interest for inducing nuclear ($\gamma,n$) reactions.
Figure 3: Measured boron doping profile. A 250 x 250 µm² crater allowed for the analysis of the material as a function of depth with ToF-SIMS. On the left scale the boron concentration is shown, on the right one the corresponding relative lattice expansion $\Delta$a/a, according to the "atomic volume interpolation" [10], in the following called modified Vergard’s law.
Figure 4: Depth profile taken with the Rocking Curve Imaging (RCI) technique [11, 12]. The signal is the integrated intensity of Bragg diffraction peaks at (111) planes with 17 keV photons.
Results
Superlattice
[…]
The boron doping profile was examined with the Time-of-Flight Secondary Ion Mass Spectrometry (ToF-SIMS) method. Figure 3 shows the depth profile. The SIMS data indicate that the growth process reproduce the four sinusoidal periods, irrespective of the analyzed region. As designed, the layers are consistent with a period length of  3.54 $\mu$m along the [100] growth direction. A slight variation of the doping maxima and minima from one period to the other can be observed for the layers apart from the substrate.
Fig. 4 depicts an enlarged map of the integrated intensity of the Bragg peak, taken with the Rocking Curve Imaging (RCI) technique [11, 12] at the European Synchrotron Radiation Facility, Grenoble. The depth profile in horizontal direction is shown by the vertical strips. The structure corresponds to the doping profile of Fig. 3 when interpreting low intensity (dark blue) with the region of large boron doping. In vertical direction the image shows that the layers are very homogeneous. This result has been confirmed over more than 1 mm along the crystal. A detailed quantitative analysis is in progress.
Figure 5: Experimental setup at MAMI, not to scale. Radiation can be observed at virtual collinear geometry of the electron beam direction and the sodium iodide NaI(Tl) radiation detector. The inset depicts the undulator crystal with the position of three SIMS craters.
Accelerator configuration
The experimental setup at the Mainz Microtron MAMI is shown at Fig. 5. The electron beam is focused onto the diamond crystal in the target chamber with a goniometer, and thereafter deflected by a dipole magnet into the beam dump. The target can be rotated in all 3 spatial directions with an accuracy of 35 $\mu$rad. The target is aligned using the signal of the ionization chamber that registers the scattered electrons. The orientation of the electron beam parallel to planes increases the scattering probability and, in turn, the signal of the ionization chamber. This method allows rapid alignment to a crystal plane.
Figure 6: (a) Experimental spectra of the undulator chip per electron, including the 194 $\mu$m thick backing crystal. Spectra taken with the NaI(Tl) detector. Spectrum in green colour for de-tuning the aperture from forward direction by $\Delta\theta_{x}$ = - 0.2 mrad, and in blue colour by + 0.2 mrad. Spectra taken at typical beam current of 3.95 pA. (b) Low energy part of difference spectrum. Background spectrum taken at $\theta_{x}$ = +0.2 mrad aperture position (blue) subtracted from spectrum taken at -0.2 mrad (green). (c) and (d) Monte-Carlo simulated spectra per electron without a backing crystal. Ideal sinusoidal boron doping profile and infinitesimal small aperture size assumed. Mean of 200 single simulations.
Spectra
The peak search was performed at an electron energy of 855 MeV. The angle of the tungsten aperture was varied around the forward direction. There exist two favourable observation angles. One is the direct forward direction $\theta_{x}=0$ for which the radiation direction defining aperture is on-axis with the electron beam direction, i.e., in Fig. 1 with the $z$ axis. The other one is the undulator radiation direction in which the aperture is de-tuned by the angle $\theta_{x}=\delta$ = - 0.174 mrad into the $z^{\prime}$ direction. For both directions $\theta_{y}=0$ was chosen.
The peak energy of the undulator radiation to be sought for can be estimated with the well known relation
$\hbar\omega=k\frac{4\pi\gamma^{2}\hbar c}{\lambda\left(1+K^{2}/2+\gamma^{2}(\delta-\theta_{x})^{2}\right)}$ (1)
with Lorentz factor $\gamma$ = 1674.2, the reduced Planck constant $\hbar$, and $c$ the speed of light. In lowest order $k$ = 1 a peak energy $\hbar\omega$ = 1.33 MeV results at observation in undulator forward direction $\theta_{x}$ = $\delta$, $A_{U}$ = 0.138 nm, $\lambda$ = 5.0 $\mu$m, and undulator parameter $K=\gamma\cdot A_{U}\cdot 2\pi/\lambda$ = 0.291.
Discussion
Spectra
The spectrum in green color of Fig. 6 exhibits a clear narrow peak with an energy of 1.30 MeV. It appears only at a negative observation angle $\theta_{x}$ = - 0.2 mrad, while at the positive angle of + 0.2 mrad mainly background is observed, see spectrum in blue colour. The result shows that our method for a peak search, as proposed in [7, Fig. 11] and applied for an unsuccessful experiment, is rather sensitive and enables the creation of a difference spectrum shown in Fig. 6 (b). In contrast to the above mentioned Si1-xGex strained layer crystal, the diamond undulator structure cannot be separated from the host diamond crystal. As a consequence, one is always faced with radiation from the backing, the thickness of which can be minimized to values in the several $\mu$m range. However, in our experiment the thickness in [100] direction was 194 $\mu$m, an actually rather large value in comparison with the undulator layer - actually a factor of 14 thicker. Therefore, a huge channeling radiation contribution is present, and consequently the difference spectrum is afflicted with large uncertainties, in particular above about 2 MeV.
Figs. 6 (c) and (d) depict Monte-Carlo simulated spectra under idealized conditions. The calculations are based on the formalism of [7], extended to simulate also sinusoidal profiles. Details are described in Methods Monte-Carlo simulation calculations. With some care in mind, the simulated spectrum (d) will be compared with the experimental one (b). The simulation reproduces with 1.29 MeV well the measured peak energy of 1.30 MeV. However, the experimental width is broader, and the total intensity significantly lower. The Bragg diffraction image of Fig. 4 may indicate that the amplitude decreases by relaxation of the stress, a fact not taken into account in the simulation calculation. In addition, an infinitesimal small aperture was assumed which is also unrealistic, not to mention that also the simulation model has systematic errors.
Finally we mention that the peak should also be observable by aligning the crystal into the (111) plane for which an enhanced peak signal is expected, since the transverse window in which the electron is captured for channeling may increase in comparison with the (110) alignment.
Prospects
Based on our observation in combination with the Monte-Carlo simulation result, we address the question of applications of narrow-band $\gamma$ radiation emitted from a boron-doped diamond superlattice. Main restrictions originate from the superlattice structure, rather than the electron accelerator facilities for which technologies are well developed. The basic underlying sensitive parameter is the relaxation length $L_{r}$ of the undulating structure as function of period length $\lambda$, number of periods $N_{U}$, amplitude $A_{U}$ and doping profile. It is probable a rather involved quest of material science which must be tackled experimentally, perhaps assisted by simulation calculations.
In the following we make the idealized assumption of an undulator with $L_{r}=\infty$, $\lambda$ = 5.0 $\mu$m, $N_{U}$ = 8, and $A_{U}$ = 0.092 nm. The boron doping profile was assumed to be triangular with a maximum boron concentration $c_{b,max}$ = 6.0 $\cdot 10^{20}$/cm3 resulting in a constant radius of curvature of 10.8 mm. In contrast to a sinusoidal doping profile there are no regions with infinite bending radius and vanishing acceleration. As a consequence, the radiation emission is maximized. An electron beam with an energy of 3 GeV was assumed. These parameters were identified from Monte-Carlo simulation calculations to produce a secondary narrow-band $\gamma$ radiation beam well suited for applications with photonuclear ($\gamma$,n) processes. The chosen parameter set does not originate from any optimization processes.
Figure 7: Simulated photon spectrum per electron. Beam energy 3 GeV, assumption of an ideal triangular boron doping profile, parameters as specified in the text. An infinitesimal small aperture was assumed, positioned at $\theta_{x}=\delta=-0.116$ mrad, and $\theta_{y}=0$. The narrow structures at 30 MeV and 46 MeV are interpreted as the second and third harmonic of the dominating peak at 14.3 MeV. A mean of 200 single simulations is shown.
[…]
@@ Discussion
The following crude estimates will be performed for the medical important 100Mo($\gamma$,n)99Mo reaction, as an example, with 10 enriched 100Mo disks of 3 mm diameter and 1 mm thickness each as secondary target, one after another. The total mass is 0.7 g. With the reaction cross section of 0.15 b [2, Fig. 4] and an irradiation time to reach half the saturation activity one obtains a specific activity of about 4.5 GBq of 99Mo (T1/2 = 65.9 h).
@@
Although the proposed generator with the assumed parameters may be less intense in comparison with existing and proposed ones [13, 14], a detailed feasibility study with optimized undulator parameters and electron beam energy seems to be worth the effort. It should include considerations on the high energy channeling background, the production of other medical interesting radionuclides, as well as considerations on the radiation damage of the diamond undulator crystal.
Methods
Superlattice design and fabrication
The design profile of the boron doping in [100] direction is of sinusoidal shape with the minimum and maximum boron content $x_{b,min}$ = 5.67 $\cdot 10^{-4}$ and $x_{b,max}$ = 56.7 $\cdot 10^{-4}$, corresponding to a boron concentration $c_{b,min}$ = 1.0 $\cdot 10^{20}$/cm3 and $c_{b,max}$ = 10.0 $\cdot 10^{20}$/cm3, respectively. Notice, a boron content is defined here as $x_{b}=n_{b}/n_{C}$ with $n_{b}$ the boron concentration in units cm-3 and $n_{C}=1.763\cdot 10^{23}$/cm3 the density of carbon atoms for diamond. The boron content reads as function of the $z$ coordinate for the first half period $x_{b}(z)=x_{b,min}+(x_{b,max}-x_{b,min})/2\cdot(1-\cos[2\pi z/\lambda])$, $20\leq z\leq\lambda/$ 2. The continuation to $\lambda/2<z\leq\lambda$ can be obtained by symmetry considerations.
The lattice expansion for the "atomic volume interpolation" [10, Fig. 1], also called linear "Modified Vegard’s low" reads $\Delta d_{110}/d_{110}(z)=k\cdot x_{b}$ with the slope constant $k=0.0962$. A differential equation $dx_{b}/dz=-k\cdot x_{b}(z)$ must be solved to obtain the nominal bending of the (110) channel. Notice that beside the boundary condition $x_{b}(z\rightarrow 0)=0$ also $dx_{b}/dz(z\rightarrow 0)=0$ must be fulfilled. The solution is schematically shown in Fig. 1.
Monte-Carlo simulation calculations
We recall from Ref. [7, Eq. (6) and (11)] that the spectral power density can be calculated if the angular trajectories $\vartheta_{x,y}(z)$ are known. The simulation calculation code described in that paper holds only for a triangular doping profile resulting in a constant bending radius for each half period. However, as described in Methods Superlattice design and fabrication the superlattice was produced with a sinusoidal doping profile. This fact required some modifications of the simulation code.
Let $u(x^{\prime})$ be the planar potential in the $(x^{\prime},z^{\prime})$ coordinate system, see Fig. 1. Than a family of sinusoidal varying potentials may be described by
with $d_{p}$ the planar distance. Transformation into the $(x,z)$ system by rotation with an angle $\delta$, taking into account the centrifugal potential, gives
$U(x,z)=u(\overline{x})-pv/R(z)~{}x$ (3)
$1/R(z)=(1/R_{min})\sin(2\pi z/\lambda)$ (5)
. In Eq. (4 )was approximated $\overline{z}=z$, $x^{\prime}=x$, $\sin(\delta)=\delta$, and $\cos(\delta)=1$. It is $pv$ = $\beta\gamma m_{e}c^{2}$ with $m_{e}c^{2}$ the rest mass of the electron and $\beta=v/c=\sqrt{1-1/\gamma^{2}}$ the reduced speed $v$ of the electron. The additional term describes the centrifugal potential with $R(z)$ the bending radius, which is a function of the longitudinal coordinate $z$, and $R_{min}$ its minimum value. The two functions $u(\overline{x})$ and $-(pv/R(z))~{}x$ are depicted in Fig. 2 and Fig. 8, upper part, respectively. The superimposed potential is shown in Fig. 8, lower part.
Figure 8: Upper: Centrifugal potential part $-(pv/R(z))~{}x$ of Eq. (3). As the particle proceeds in $z$ direction the potential part $u(\overline{x})$ depicted in Fig. 2 is modified for any specific $z$ by the centrifugal term. Only the positive part of the potential is shown. Lower: Total transverse potential landscape $U(x,z)$, Eq. (3). The big hills at right and left originate from the centrifugal potential. However, gradient forces originating from $-(pv/R(z))~{}x$, see upper part, act only in $x$ direction.
An electron entering the potential landscape at the window close to $x=z$ = 0 may be captured in the shown channel. Collisions may result in de-channeling and the free particle with a positive transverse energy may bounce against the right potential hill, will be reflected and may enter in regions of negative centrifugal potential below the plane $U(x,z)$ = 0. As it proceeds further in $z$ direction it may be reflected by an increasing centrifugal potential, like the left hill, in opposite direction.
[…]
For each $\Delta z_{k}$ the scattering angle changes $\Delta\vartheta_{k}^{at}$ and $\Delta\vartheta_{k}^{el}$ by collisions with atoms and electrons, respectively, are determined randomly by a Monte Carlo simulation. The new scattering angle is given with $\Delta\vartheta_{k}^{sc}$ = $\Delta\vartheta_{k}^{at}$ + $\Delta\vartheta_{k}^{el}$ by
$\vartheta_{k+1}(z_{k})$ $=$ $\vartheta_{k}^{kin}+\Delta\vartheta_{k}^{sc}=$ (8)
=±$=\pm\sqrt{2\big{(}E_{\bot,k}-U(x_{k},z_{k})\big{)}/pv}+\Delta\vartheta_{k}^{sc}.$
The first term $\vartheta_{k}^{kin}$ is the instantaneous scattering angle associated with the kinetic energy $E_{\bot,k}-U(x_{k},z_{k})$. Notice, at move (ii) forces are assumed to be absent, meaning that the kinetic energy $T_{\bot,k+1}$ remains constant and the scattering angle does not change.
$E_{\bot,k+1}-E_{\bot,k}$ $=$ (9)

### v2

# Observation of narrow-band $\gamma$ radiation from a boron-doped diamond superlattice with an 855 MeV electron beam

@@ Abstract
I Introduction
Intense narrow-band multi-MeV photon beams have a large potential in fields of basic and nuclear research, medicine and technical applications [1, 2, 3, 4, 5]. Having in mind in particular photonuclear reactions for application purposes, the production mechanisms of such beams should be targeted, from an economical point of view, on electron accelerators with moderate energies in the GeV range or below. Bremsstrahlung radiation created by an electron beam incident on a high-Z target has been widely used for many decades. However, such sources features broad band radiation characteristics. Nearly mono-energetic $\gamma$-ray beams of high intensity can be produced by facilities based on laser Compton scattering which are already operational [6] or currently under construction [4, Table 10].
Figure 1: Transverse potential $U(x,z)$ for a single (110) channel of a boron doped diamond superlattice. Shown is one period with length of $\lambda$ = 5.0 $\mu$m. Coordinates $(x,z)$ of a full sinusoidal period (0,0) and (-8.7 Å, 5 $\mu$m) define the $z^{\prime}$ undulator direction around the potential wiggles. It makes an angle $\delta$ = -0.174 mrad with $z$ axis, the nominal electron beam direction. Notice, longitudinal $z$ coordinate is squeezed by four orders of magnitudes. For details see appendix VII.2. Barrier heights are modified as functions of $z$ by centrifugal forces. Insets show examples, $x_{1}$ = -0.78 Å and $x_{2}$ = -7.90 Å. Figure based on design parameters of section II.1.
In this work an alternative route towards narrow band photon sources was elaborated which is based on the channeling process of ultra-relativistic electrons in periodical bent crystals. The method, theoretically known since long, has been comprehensively treated in monographs [7, 8] in which all relevant references can be found. It resembles magnetic undulators, however, with period lengths in the multi-$\mu$m range rather than on a centimeter scale. The principle will be explained by means of Fig. 1. We use the continuum potential picture introduced by Lindhard [9], applied also in our case-study [10] on which this publication is based.
An electron entering the crystal at the coordinates $x=z$ = 0 nearly parallel to the $z$ axis may be captured in the potential pocket, which at that point is symmetrical with a barrier hight of 24 eV It performs oscillations, thereby following the bent sinusoidal shape of the (110) plane. By means of the accompanied acceleration channeling radiation and undulator-like radiation will be emitted. The radiation is strongly forward directed, for undulator-like radiation typically in the order of 0.5 mrad for 1 GeV electrons. The emitted photon energy is in the $\gamma$-ray domain, for a period length of a few $\mu$m typically in the MeV range. It scales with the electron beam energy squared and inversely with the undulator period length $\lambda$. The capture probability and in turn the intensity of the $\gamma$ radiation requires high quality electron beams with a low angular spread. The reason is that the transverse energy $pv/2\cdot\vartheta_{x}^{2}$ is the initial energy of the particle in the potential pocket with $pv$ in essence the electron beam energy and $\vartheta_{x}$ the angle of the electron with respect to the $z$ axis.
The experimental challenge is to fabricate crystalline undulators with such small period lengths. A graded composition strained layer in a superlattice was proposed [11]. Experiments were performed with Si1-xGex undulator crystals grown at the Aarhus University [12, 13], however, results were not such convincing that applications discussed in [14] were pursued.
An alternative are diamond undulators which can be produced by Chemical Vapour Deposition (CVD) with periodically varying content of boron. The effect is based on the functional dependence of the lattice constant on the boron content, see e.g. [15, Fig. 1]. The low $Z$ = 6 material diamond has the advantage, beside its radiation hardness, that scattering of the beam electrons by atomic nuclei and electrons comprising the crystal is reduced in comparison with silicon, $Z$ = 14, which potentially results in an increase of the transverse energy and de-channeling processes. This is the case, in particular, in regions in which the radius of curvature is small, i.e., centrifugal forces are large and one of the barrier heights is reduced, cf. Fig. 1 at $z_{1}$ = 1.25 $\mu$m and $z_{2}$ = 3.75 $\mu$m. - Unsuccessful experiments were described in [16, 10].
Crucial is, beside the fabrication of a high quality superlattice by the CVD process, a strategy for the peak search at the expected energy which was developed in a preceding case study [10]. To this end Monte-Carlo simulation calculations were performed with a computer code developed by us, which fairly well reproduces the basic measured characteristics of the quasi-monochromatic $\gamma$ radiation as the peak energy, its width and the intensity. This fact encouraged us to predict, as an example, features of intense quasi-monochromatic 14.3 MeV radiation from an eight-period diamond undulator with $\lambda$ = 5 $\mu$m and a 3 GeV electron beam. Such a $\gamma$-ray source may be of potential interest for inducing nuclear ($\gamma,n$) reactions.
The paper is organized as follows. In section II the superlattice is described, including diagnosis of it, and the experimental setup at MAMI. Section III is devoted to the peak search. The result is discussed in section IV by means of a comparison with Monte-Carlo simulation calculations. In the follows section V possible future applications are sketched on the example of the production of the medical important 99Mo/99mTc isotopes. The paper closes in section VI with conclusions. Some details of the superlattice preparation as well as of the Monte-Carlo simulation calculations are moved into the appendices VII A and B, respectively.
II Experimental
II.1 The diamond undulator
Parameters of the diamond superlattice were optimized by means of Monte-Carlo simulation calculations for an experiment with an electron beam energy of 855 MeV. The period length must be chosen such that the undulator peak energy is well separated from the channeling radiation distribution. Finally we found the following design parameters: period length in [100] direction 3.54 $\mu$m, number of periods $N_{U}$ = 4, and amplitude $A_{U}$ = 0.138 nm. Details are described in appendix VII.1.
Figure 2: (a) Measured boron doping profile. A 250 x 250 µm² crater allowed for the analysis of the material as a function of depth with ToF-SIMS. On the left scale the boron concentration is shown, on the right one the corresponding relative lattice expansion $\Delta$a/a, according to the ”atomic volume interpolation” [15], also called ”modified Vegard’s law”. (b) Depth profile obtained with the Rocking Curve Imaging (RCI) technique [17, 18]. The signal is the integrated intensity of Bragg diffraction peaks at (111) planes with 17 keV photons.
The boron doping profile was examined with the Time-of-Flight Secondary Ion Mass Spectrometry (ToF-SIMS) method. Fig. 2 (a) shows the depth profile. The ToF-SIMS data indicate that the growth process reproduce the four sinusoidal periods, irrespective of the analyzed region. As designed, the layers are consistent with a period length of  3.54 $\mu$m along the [100] growth direction. A slight variation of the doping maxima and minima from one period to the other can be observed for the layers apart from the substrate.
Fig. 2 (b) depicts a section of an integrated intensity map of the Bragg peak, taken with the Rocking Curve Imaging (RCI) technique [17, 18, 19] at the European Synchrotron Radiation Facility, Grenoble. The structure corresponds to the doping profile of Fig. 2 (a). This result has been confirmed over more than 1 mm along the crystal.
Figure 3: Experimental setup at MAMI, not to scale. Radiation can be observed at virtual collinear geometry of the electron beam direction and the sodium iodide NaI(Tl) radiation detector. The inset depicts the undulator crystal with the position of three ToF-SIMS craters.
II.2 Experimental setup at MAMI
The experimental setup at the Mainz Microtron MAMI is shown at Fig. 3. The electron beam is focused onto the diamond crystal in the target chamber with a goniometer, and thereafter deflected by a dipole magnet into the beam dump. The target can be rotated in all 3 spatial directions with an accuracy of 35 $\mu$rad. The target is aligned using the signal of the ionization chamber that registers scattered electrons. The orientation of the electron beam parallel to planes increases the scattering probability and, in turn, the signal of the ionization chamber. This method allows rapid alignment to a crystal plane.
Figure 4: (a) Experimental spectra of the undulator chip per electron, including the 194 $\mu$m thick backing crystal. Spectra taken with the NaI(Tl) detector at a typical beam current of 3.95 pA. Spectrum in green colour for de-tuning the aperture from $z$ direction by $\theta_{x}$ = - 0.2 mrad, and in blue colour by + 0.2 mrad. (b) Low energy part of difference spectrum. Background spectrum taken at $\theta_{x}$ = +0.2 mrad aperture position (blue) subtracted from spectrum taken at -0.2 mrad (green). (c) and (d) Monte-Carlo simulated spectra per electron without a backing crystal. Ideal sinusoidal boron doping profile and infinitesimal small aperture size assumed. Mean of 200 single simulations.
III Measurements and results
The peak search was performed at an electron energy of 855 MeV. The angle of the tungsten aperture was varied. There exist two favourable observation angles. One is the direct forward direction $\theta_{x}=0$ for which the radiation direction defining aperture is on-axis with the electron beam direction. It is in Fig. 1 the $z$ axis which coincides with the [110] direction of the host crystal. The other one is the undulator direction in which the aperture is de-tuned by the angle $\theta_{x}=\delta$ = - 0.174 mrad into the $z^{\prime}$ direction. For both directions $\theta_{y}=0$ was chosen. For details see appendix VII.1.
The peak energy of the undulator radiation to be sought for can be estimated with the well known undulator relation [10, Eq. (7)] to be 1.33 MeV at observation in undulator $z^{\prime}$ direction $\theta_{x}$ $\delta$ = 0, $A_{U}$ = 0.138 nm, $\lambda$ = 5.0 $\mu$m, undulator parameter $K=\gamma\cdot A_{U}\cdot 2\pi/\lambda$ = 0.291, and $\gamma$ = 1674.2 for 855 MeV.
IV Discussion
The spectrum in green color of Fig. 4 (a) exhibits a clear narrow peak with an energy of 1.30 MeV. It appears only at a negative observation angle $\theta_{x}$ = - 0.2 mrad, while at the positive angle of + 0.2 mrad mainly background is observed, see spectrum in blue colour. The result shows that our method for a peak search, as proposed in [10, Fig. 11] and applied for an unsuccessful experiment, is rather sensitive and enables the creation of a difference spectrum shown in Fig. 4 (b). This crucial methodical procedure was not addressed in [20, 21]. In contrast to the above mentioned Si1-xGex strained layer crystal, the diamond undulator structure cannot be separated from the host diamond crystal. As a consequence, one is always faced with radiation from the backing, the thickness of which can be minimized to values in the several $\mu$m range. However, in our experiment the thickness in [100] direction was 194 $\mu$m, a rather large value in comparison with the undulator layer thickness of 3.54 $\mu$m. Therefore, a huge channeling radiation contribution is present, although somehow suppressed due to the choice of the observation direction, and consequently the difference spectrum is afflicted with large uncertainties, in particular about 2 MeV.
Figs. 4 (c) and (d) depict Monte-Carlo simulated spectra under idealized conditions. The calculations are based on the formalism of [10], extended to simulate also sinusoidal profiles. Details are described in appendix VII.2. The bremsstrahlung contribution was estimated from [10, Fig. 9 (c)] to be small and has been neglected. With some care in mind, the simulated spectrum (d) will be compared with the experimental one (b). The simulation reproduces with 1.29 MeV well the measured peak energy of 1.30 MeV. However, the experimental width is broader, and the total intensity significantly lower. The Bragg diffraction image of Fig. 2 (b) may indicate that the amplitude decreases by relaxation of the stress, probably my misfit dislocations [22], a fact not taken into account in the simulation calculation. In addition, an infinitesimal small aperture was assumed which is also unrealistic, not to mention that also the simulation model has systematic errors.
Finally we mention that a peak should also be observable by aligning the crystal into the (111) plane.
V Prospects
Based on our observation in combination with the Monte-Carlo simulation result, we address the question of applications of narrow-band $\gamma$ radiation emitted from a boron-doped diamond superlattice. Main restrictions originate from the superlattice structure, rather than the electron accelerator facilities for which technologies are well developed. The basic underlying sensitive parameter is the relaxation length $L_{r}$ of the undulating structure as function of period length $\lambda$, number of periods $N_{U}$, amplitude $A_{U}$ and doping profile. It is probable a rather involved quest of material science which must be tackled experimentally, perhaps assisted by realistic Monte-Carlo simulation calculations of the type described in [23].
In the following we make the idealized assumption of an undulator with $L_{r}=\infty$, $\lambda$ = 5.0 $\mu$m, $N_{U}$ = 8, and $A_{U}$ = 0.092 nm. The boron doping profile was assumed to be of triangular shape with a maximum boron concentration $c_{b,max}$ = 6.0 $\cdot 10^{20}$/cm3 resulting in a constant radius of curvature of 10.8 mm. In contrast to a sinusoidal doping profile there are no regions with infinite bending radius and vanishing acceleration. As a consequence, the radiation emission is maximized. An electron beam with an energy of 3 GeV was assumed. These parameters were identified from Monte-Carlo simulation calculations to produce a secondary narrow-band $\gamma$ radiation beam well suited for applications with photonuclear ($\gamma$,n) processes. The chosen parameter set does not originate from any optimization processes.
Figure 5: Simulated photon spectrum per electron. Beam energy 3 GeV, assumption of an ideal triangular boron doping profile, parameters as specified in the text. An infinitesimal small aperture was assumed, positioned at $\theta_{x}=\delta=-0.116$ mrad, and $\theta_{y}=0$. The narrow structures at 30 MeV and 46 MeV are interpreted as the second and third harmonic of the dominating peak at 14.3 MeV. Mean of 200 single simulations.
@@ V Prospects
The following crude estimates will be performed for the medical important 100Mo($\gamma$,n)99Mo reaction, as an example, with 10 enriched 100Mo disks of 3 mm diameter and 1 mm thickness each as secondary target, one after another. The total mass is 0.7 g. With the reaction cross section of 0.15 b [3, Fig. 4] and an irradiation time to reach half the saturation activity one obtains a specific activity of about 4.5 GBq of 99Mo (T1/2 = 65.9 h).
@@
Although the proposed generator with the assumed parameters may be less intense in comparison with existing and proposed ones [24, 25], a detailed feasibility study with optimized undulator parameters and electron beam energy seems to be worth the effort. It should include considerations on the high energy channeling background, the production of other medical interesting radionuclides or applications in other fields, as well as considerations on the radiation damage of the diamond undulator crystal.
VI Conclusion
A 4-period diamond superlattice was produced by a CVD process with a sinusoidal boron doping profile grown on a (100) diamond host crystal of high quality. The accompanied periodic variation of the lattice constant resulted in sinusoidally deformed (110) planes with a period length of 5.0 $\mu$m and an amplitude of 0.138 nm. A channeling experiment was performed with the 855 MeV electron beam of the Mainz Microtron MAMI accelerator facility. A clear peak was observed with a large sodium iodide scintillation detector at a photon energy of 1.30 MeV. Monte-Carlo simulation calculations, assuming an ideal sinusoidal undulator profile, reproduce well the photon peak energy. However, the measured line width and intensity seem, in comparison, to be broadened and reduced, respectively, for which explanations have been discussed.
Prospects were outlined by the features of an 14.3 MeV intense $\gamma$-ray beam as modelled by simulation calculations. Potential applications with rather small target samples are conceivable in fields like basic nuclear research, technical applications or even medicine. The latter was exemplary examined by the photonuclear 100Mo($\gamma$,n)99Mo reaction for the production of the medical important 99mTc isotope.
Data availability statement
All data are freely available.